A battery, capacitor, resistor, and switch are con

How do I do all parts above?
Answer #1

a) just after potential across the capacitor =0

and potential across resistor =V (emf of battery =V)

b) charge inside the capacitor =

V =iR+Q/C (Q =di/dt)

so Q =CV(1- e-t/RC)

d) so potential difference across capacitor = Q/C=V(1- e-t/RC)

electric field inside the capacitor =v/d =V(1- e-t/RC)/d

c) force =QE = CV(1- e-t/RC)*V(1- e-t/RC)/d =CV2(1- e-t/RC)2/d